Use the RTSC approach to solve the following application.there are no existing formulas that can be used for this application.Four times the difference of a number and two is twenty-nine less than the number.find the number
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

4(n - 2) = n - 29
4n - 8 = n - 29
3n = -21

n = -7

Answer by ikleyn(52803)

.

I asked GOOGLE about the meaning of this abbreviation RTSC.

The Internet returned me back this list of possible meanings


Acronym Definition
RTSC Raytheon Technical Services Company
RTSC Return to System Call (Returning materials repaired to customer request)
RTSC Read The Source Code
RTSC Return to Sleepaway Camp (film)
RTSC Rule to Show Cause (legal)
RTSC Real-Time Signal Controller
RTSC Railway Technology Strategy Centre (UK)
RTSC Real Time Strategic Carnage (gaming website)
RTSC Real Time Supply Chain
RTSC Rural Telephone Service Company (Kansas)
RTSC Royal Thai Supreme Command
RTSC Regional Training Support Center
RTSC Real-Time Software Component
RTSC Recreational Training Scuba Council
RTSC Room Temperature Superconductor
RTSC Responds to Simple Commands
RTSC Regional Technology Service Center(s) (Sprint)
RTSC Remuneration and Terms of Service Committee
RTSC Regional Training Simulation Center
RTSC Reeds Target Shooting Club (UK)
RTSC Real-Time System Component (TI Technology)


Which one of these I am supposed to use ?


////////////////

Totally idiotic instruction.